OpenStudy (anonymous):

I Need Help Asap ... Giving Testimonial Medal And Fan !! Find f(5) for this sequence: f(1) = 2 and f(2) = 3, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______

OpenStudy (anonymous):

F(5) = 22

OpenStudy (giobott09):

f(5) = 6+16 f(5) = 22

OpenStudy (anonymous):

thanks so much :)

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

sorry i cant give you a medal but i already became your fan @yowzaman

OpenStudy (anonymous):

No worries :)

OpenStudy (giobott09):

@yowzaman Medal for medal?

OpenStudy (anonymous):

Can you help me w/ another one @yowzaman Generate the first 5 terms of this sequence: f(1) = 2 and f(2) = 3, f(n) = f(n - 1) + f(n - 2), for n > 2. A) 2, 3, 5, 7, 9 B) 2, 3, 4, 7, 11 C) 2, 3, 5, 9, 11 D) 2, 3, 5, 8, 13

OpenStudy (anonymous):

@Giobott09 sure :)

OpenStudy (anonymous):

I think the answer is D

OpenStudy (anonymous):

Okay thanks :)